OpenStudy (poofypenguin):

What is the absolute maximum value taken by the function f(x) = 5sin(x)-2cos(x) on [0, pi/2]?

hartnn (hartnn):

are u stuck at 5cos x +2 sin x =0 ?

OpenStudy (poofypenguin):

yeah... i'm trying to find the critical values again... but i can't seem to find any... does that mean there arn't any?

hartnn (hartnn):

i think there are... divide the entire equation by \(\sqrt{5^2+2^2}\) then put 5/sqrt[...] as sin a and 2/ sqrt[..] as cos a

hartnn (hartnn):

and u get the form sin (x+a) =0

hartnn (hartnn):

getting this ?

OpenStudy (poofypenguin):

not really... is it possible you could work it out so i can see how it works?

hartnn (hartnn):

yeah..will take some time...

hartnn (hartnn):

\(5cos x +2 sin x =0\) \(\huge \frac{5cosx}{\sqrt{5^2+2^2}}+\frac{2sin x}{\sqrt{5^2+2^2}}=0\) put sin a = 5/sqrt[29], then 2/sqrt[29] will be cos a \(\sin a\cos x+\cos a\sin x=0 \\ sin[x+a]=0\) ok till here ?

hartnn (hartnn):

where a is \(\tan^{-1}(5/2)\) or \(\sin^{-1}(5/\sqrt{29})\)

OpenStudy (poofypenguin):

I understand the first part, but where did the a's come from?

hartnn (hartnn):

a is just a constant. i denoted 5/sqrt 29 as sin a to get the form sin a cos b +cos a sin b to write that as sin(a+b)

OpenStudy (poofypenguin):

ok...i think i get it so far...

hartnn (hartnn):

sin (x+a) =0 so x+a =.... ?

hartnn (hartnn):

sin(x+a) = 0 since, sin 0 = 0 u get x+a = 0 x=-a now since a is positive, x is negative, which is not in the range of [0,pi/2] so again to find absolute maxima, substitute end-points put x=0 and x=pi/2 whichever is greater is your maxima.

OpenStudy (poofypenguin):

Alright! I think i pretty much get what you're saying.... i just need to stare at it a little yet! Thanks so much again! :D

hartnn (hartnn):

welcome ^_^